# Ctrl-GenAug: Controllable Generative Augmentation for Medical Sequence Classification

### v2

@@ 3.3 Multimodal Conditions Guidance
3.3.1 Semantic Conditions for Visual Appearance Control
As shown in Fig. 4, the sequence generator exploits three semantic conditions to perform visual appearance control, thus achieving controllable and high-fidelity synthesis.
- Descriptive Text: It provides an intuitive indication of the coarse-grained semantic concepts of sequences. In our implementation, we pre-align the visual and textual features by finetuning the widely used CLIP model (Radford et al., 2021) using our medical datasets to mitigate the large gap between them. Then, semantic information of texts can feasibly query relevant medical visual patterns, thus easing subsequent training. Similar to (Rombach et al., 2022), the text guides the generator via cross-attention layers.
@@ 3.3.1 Semantic Conditions for Visual Appearance Control
[…]
@@
3.3.2 Sequential Condition for Serial Guidance
Apart from perceiving visual concepts, modeling sequential knowledge is also important in sequence synthesis, with sequential cues playing a vital role. Most existing methods (Liang et al., 2023; Ni et al., 2023) adopted dense optical flow (Zach et al., 2007) to promote sequential dynamics modeling. However, optical flow extraction requires high computational demands (Wang et al., 2024b). Thus, we instead introduce computation-efficient motion field as a sequential condition in sequence LDM. It explicitly showcases the pixel-wise motions between adjacent frames/slices (Figs. 3-4). In our approach, we first extract the motion fields of real sequences using a Python package (Bommes et al., 2020). Then, a motion encoder (Fig. 4) receives the motion fields and produces motion features. Last, the features are concatenated with latent representations $\mathit{z}$ from VAE along the channel dimension for serial guidance.
@@ 3.4 Sequential Augmentation Module
3.4.1 Key-frame/slice Attention
The common spatial self-attention layers can lead to sequential inconsistency due to the lack of interaction across frames/slices. To augment sequential consistency, we introduce a key-frame/slice attention (KA) mechanism, where two selected key-frames/slices act as references to propagate S2 information throughout the sequence. Specifically, for any frame/slice, we select its previous and the first counterpart of the sequence as the references and transform spatial self-attention into KA, aiming to align the latent features $z_{l}$ of the l-th frame/slice with $z_{1}$ and $z_{l-1}$. We obtain query from $z_{l}$, key and value features from $z_{1}$ and $z_{l-1}$, and compute $Attention\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)$ using:
@@ 3.4.1 Key-frame/slice Attention
[…]
3.4.2 Motion Field Attention
To further boost the sequential coherence of the generated sequences, we propose to reuse the motion field condition (refer to Sec. 3.3.2) by introducing a motion field attention (MFA) mechanism after KA. As shown in Fig. 5(c), MFA requires the patches to communicate with those in the same motion field-based pathway including itself, which eliminates flickers of the generated sequences to make the contents visually smooth. Inspired by (Cong et al., 2024), MFA is implemented in two steps: a) motion field-based patch pathway sampling and b) attention calculation. In step 1, we sample the patch pathways based on the m-scaled downsampled motion fields. Take a video sequence as an example, for a patch $p_{l}$ on the l-th frame of a f-frame video, the $path$ can be derived from the motion field. Since the sampling procedure inevitably generates multiple pathways for the same patch, we randomly sample a pathway to ensure its uniqueness to which each patch belongs. In this setting, let $H,W$ the height and width of the input video frame, then the size of the pathway set after sampling equals $\frac{H\times W}{m^{2}}\times f$. In step 2, we calculate $Attention(\mathbf{Q},\mathbf{K},\mathbf{V})$ with:
@@ 3.4.2 Motion Field Attention
where $z_{path}$ represents the latent features of patches on the $path$, as follows:
@@
1:synthetic samples set $\mathrm{S}$, real data-trained classifier $p$.
@@
3:$\mathrm{S_{1}}=\varnothing$triangleright synthetic samples set obtained after stage 1
@@
6:    $l_{x}=-\sum c\mathrm{log}\sigma(p(\mathrm{S}[i][x]))$triangleright loss of the sample indexed $x$
7:    $L_{c}=\frac{1}{M}\sum_{x=1}^{M}l_{x}$triangleright noise identification threshold
8$\triangleright$ $M$ clips in the group
@@
13:             $\mathrm{S_{1}}\leftarrow\mathrm{S_{1}}\cup\{\mathrm{S}[i][x]\}$triangleright save less noisy sample
@@
17:return $\mathrm{S_{1}}$triangleright with $n_{1}(<n)$ groups totaling $N_{1}(<N)$ clips
@@
19:$\mathrm{S_{2}}=\varnothing$triangleright synthetic samples set obtained after stage 2
20:$\mathrm{A}=\left\{A_{1},...,A_{N}\right\}$triangleright VAE-Seq values set of $\mathrm{S}$
21:$\mathrm{B}=\left\{B_{1},...,B_{N_{1}}\right\}$triangleright VAE-Seq values set of $\mathrm{S_{1}}$
22:$t_{l},t_{h}\leftarrow\mathbf{KMeans}(\mathrm{A},\mathbf{K}=4)$triangleright boundary thresholds
@@
25:         $\mathrm{S_{2}}\leftarrow\mathrm{S_{2}}\cup\{y_{j}|y_{j}\in\mathrm{S_{1}}\}$triangleright save gentle dynamic sample
@@
30:return $\mathrm{S_{2}}$triangleright with $n_{2}(<n_{1})$ groups totaling $N_{2}(<N_{1})$ clips
@@
32:$\mathrm{S_{3}}=\varnothing$triangleright synthetic samples set obtained after stage 3
@@
37:             $\mathrm{S_{3}}\leftarrow\mathrm{S_{3}}\cup\mathrm{S_{2}}[k][q]$triangleright save diverse sample
@@
43:return $\mathrm{S_{3}}$triangleright final synthetic database for downstream training
@@ 3.5 Noisy Synthetic Data Filter
Although visually realistic and smooth (good cases, Fig. 6), sequence synthesis may still suffer from class semantics misalignment, cross-frame/slice inconsistency or over-consistency (i.e., almost static clip), and inter-clip similarity. For instance, in Fig. 6(e), the synthesized carotid clip category is wrong, which should be moderate rather than mild. In Fig. 6(f), the generated clip includes abrupt changes in anatomical structures. Hence, blindly using all synthetic clips for classifier learning will significantly cause a performance drop due to noisy samples. Our work proposes a noisy synthetic data filter to adaptively remove harmful generated clips at class semantics and sequential levels, as illustrated in Algorithm 1. This can also link the synthetic data to downstream tasks, thus potentially achieving a higher performance upper bound.
3.5.1 Class Semantics Misalignment Filtering
This scheme is designed to filter the generated clips whose visual contents mismatch the corresponding class labels. Motivated by anomaly detection (Kwon et al., 2020), we consider that such noisy synthetic samples will produce large losses if fed into a powerful classifier trained on real ones. Hence, a loss threshold can be set to identify and filter the anomalistic clips. Specifically, we first adopt a well-trained classifier in the real data domain to calculate case-wise cross-entropy loss values $\left\{l_{x}\right\}_{x=1}^{M}$ based on corresponding category $c$ for all synthetic clips in each group. Then, we set the average loss as the noise identification threshold of $c$. Last, we conduct class semantics misalignment filtering at the group level. If the loss of a synthetic clip with $c$ exceeds the threshold $L_{c}$, it will be viewed as a noisy sample and excluded from the synthetic pool.
3.5.2 Sequential Filtering
This strategy aims to filter noisy samples at a) inner-sequence and b) inter-sequence levels. For a), we screen out synthetic clips with gentle dynamic, avoiding those that are either inconsistent or over-consistent affecting downstream learning. Specifically, we retain a synthetic clip whose cross-frame/slice consistency falls in a pre-computed range. It is determined by K-means clustering (Macqueen, 1967) based on the cross-frame/slice consistency of all synthetic clips. To assess the video sequence consistency, (Wang et al., 2024b) tends to use CLIP (Radford et al., 2021) image embeddings to compute the average cosine similarity across consecutive frames. However, this method leads to inadequate evaluation due to the usage of limited informative CLIP embeddings. To resolve the problem, we propose a metric termed VAE-Seq that utilizes latent embeddings from pretrained VAE, instead of deriving embeddings from CLIP space, to assess cross-frame/slice consistency. Therefore, with the higher dimension of latents (e.g., 4$\times$32$\times$32 for input size 256$\times$256 compared to 768 in CLIP embeddings), VAE-Seq reflects an accurate evaluation using more fine-grained and informative features. For b), we seek to diversify the samples in the synthetic set to prevent overfitting and avoid wasting computational resources in downstream training. To this end, we perform the inter-sequence filtering based on inter-clip similarity in each group, as shown in Algorithm 1. We quantify the inter-clip similarity by calculating the cosine distance of VAE latent embeddings between frame/slice pairs in two clips and averaging the obtained distance values.
3.6 Evaluation
Existing sequence synthesis works evaluate the synthetic quality by mostly using Fréchet Inception Distance (FID) (Heusel et al., 2017) and Fréchet Video Distance (FVD) (Unterthiner et al., 2018). However, researchers have verified that these metrics do not consistently correlate with performance metrics on downstream tasks (Yang et al., 2024; Luo et al., 2024). To exhaustively evaluate the values of synthetic cases to downstream tasks, 11 popular classifiers are deployed for comparison with respect to diagnostic accuracy and area under the receiver operating characteristic curve (AUROC). Besides, we adopt the proposed VAE-Seq to assess the cross-frame/slice consistency (more details refer to Sec. 3.5.2). Moreover, motivated by (Huang et al., 2024c), we quantify the smoothness of generated clips using the metric termed Dynamic Smoothness, which is the mean absolute error between the reconstructed frames/slices from the interpolation model (Li et al., 2023) and the original real ones.
@@ 3.6 Evaluation
Figure 6: Typical good and bad synthetic samples of three datasets using different banks of conditions. Conditional controls include class label (C), text (T), image prior (I), and motion field (MF). See more synthetic results in Supplementary Material.
@@ 4 Experiments
We first comprehensively validated the effectiveness and generality of our Ctrl-GenAug on 3 medical datasets, using 11 popular networks trained on 3 paradigms. We then tested the performance of Ctrl-GenAug in underrepresented high-risk sets and out-domain conditions. Last, we analyzed the contributions of our multimodal conditions guidance, sequential augmentation module, and noisy synthetic data filter, respectively; as well as the impacts of traditional and generative augmentations.

### v3

@@ 3.3 Multimodal Conditions Guidance
Semantic Conditions for Visual Appearance Control. As shown in Fig. 4, the sequence generator exploits three semantic conditions to perform visual appearance control, thus achieving controllable and high-fidelity synthesis.
- Descriptive Text: It provides an intuitive indication of the coarse-grained semantic concepts of sequences. In our implementation, we pre-align the visual and textual features by finetuning the widely used CLIP model (Radford and others, 2021) using our medical datasets to mitigate the large gap between them. Then, semantic information of texts can feasibly query relevant medical visual patterns, thus easing subsequent training. Similar to (Rombach et al., 2022), the text guides the generator via cross-attention layers.
@@
$\left\{\begin{aligned} &\mathbf{Q_{g}}=W^{Q_{g}}f_{g},\mathbf{K_{g}}=W^{K_{g}}f_{g},\mathbf{V_{g}}=W^{V_{g}}f_{g},g\in\left\{\mathbf{t},\mathbf{i}\right\},\\ &Attention\left(\mathbf{Q_{g}},\mathbf{K_{g}},\mathbf{V_{g}}\right)=\sum_{g}\left[\sigma(\frac{\mathbf{Q_{g}K_{g}^{\top}}}{\sqrt{d}}\mathbf{V_{g}})\right],\end{aligned}\right$. (2)
@@
Sequential Condition for Serial Guidance. Apart from perceiving visual concepts, modeling sequential knowledge is also important in sequence synthesis, with sequential cues playing a vital role. Most existing methods (Liang et al., 2023; Ni et al., 2023) adopted dense optical flow (Zach et al., 2007) to promote sequential dynamics modeling. However, optical flow extraction requires high computational demands (Wang et al., 2023). Thus, we instead introduce computation-efficient motion field as a sequential condition in sequence LDM. It explicitly showcases the pixel-wise motions between adjacent frames/slices (Figs. 3-4). In our approach, we first extract the motion fields of real sequences using a Python package (Bommes et al., 2020). Then, a motion encoder (Fig. 4) receives the motion fields and produces motion features. Last, the features are concatenated with latent representations $\mathit{z}$ from VAE along the channel dimension for serial guidance.
@@ 3.4 Sequential Augmentation Module
Key-frame/slice Attention. The common spatial self-attention layers can lead to sequential inconsistency due to the lack of interaction across frames/slices. To augment sequential consistency, we introduce a key-frame/slice attention (KA) mechanism, where two selected key-frames/slices act as references to propagate S2 information throughout the sequence. Specifically, for any frame/slice, we select its previous and the first counterpart of the sequence as the references and transform spatial self-attention into KA, aiming to align the latent features $z_{l}$ of the l-th frame/slice with $z_{1}$ and $z_{l-1}$. We obtain query from $z_{l}$, key and value features from $z_{1}$ and $z_{l-1}$, and compute $Attention\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)$ using:
@@
where $W^{Q},W^{K},W^{V}$ are initialized on the original spatial self-attention weights for inheriting the semantic perception capability of LDM in the finetuning stage. $\left\{\cdot\right\}$ represents concatenation operation. It is highlighted that KA retains low computational complexity compared with full attention (Zhang et al., 2023). Please refer to Fig. 5(b) for a visual illustration.
Motion Field Attention. To further boost the sequential coherence of the generated sequences, we propose to reuse the motion field condition (refer to Sec. 3.3) by introducing a motion field attention (MFA) mechanism after KA. As shown in Fig. 5(c), MFA requires the patches to communicate with those in the same motion field-based pathway including itself, which eliminates flickers of the generated sequences to make the contents visually smooth. Inspired by (Cong et al., 2024), MFA is implemented in two steps: a) motion field-based patch pathway sampling and b) attention calculation. In step 1, we sample the patch pathways based on the m-scaled downsampled motion fields. Take a video sequence as an example, for a patch $p_{l}$ on the l-th frame of a f-frame video, the pathpath can be derived from the motion field. Since the sampling procedure inevitably generates multiple pathways for the same patch, we randomly sample a pathway to ensure its uniqueness to which each patch belongs. In this setting, let $H,W$ the height and width of the input video frame, then the size of the pathway set after sampling equals $\frac{H\times W}{m^{2}}\times f$. In step 2, we calculate $Attention(\mathbf{Q},\mathbf{K},\mathbf{V})$ with:
@@
where $z_{path}$ represents the latent features of patches on the pathpath, as follows:
@@
1:synthetic samples set $\mathrm{S}$, real data-trained classifier pp.
@@
[…]
@@
6:  $l_{x}=-\sum c\mathrm{log}\sigma(p(\mathrm{S}[i][x]))$ $\triangleright$ loss of the sample indexed $x$
7:  $L_{c}=\frac{1}{M}\sum_{x=1}^{M}l_{x}$ $\triangleright$ noise identification threshold
[…]
@@
13:     $\mathrm{S_{1}}\leftarrow\mathrm{S_{1}}\cup\{\mathrm{S}[i][x]\}$ $\triangleright$ save less noisy sample
@@
17:return $\mathrm{S_{1}}$ $\triangleright$ with $n_{1}(<n)$ groups totaling $N_{1}(<N)$ clips
@@
[…]
@@
25:    $\mathrm{S_{2}}\leftarrow\mathrm{S_{2}}\cup\{y_{j}|y_{j}\in\mathrm{S_{1}}\}$ $\triangleright$ save gentle dynamic sample
@@
30:return $\mathrm{S_{2}}$ $\triangleright$ with $n_{2}(<n_{1})$ groups totaling $N_{2}(<N_{1})$ clips
@@
[…]
@@
37:     $\mathrm{S_{3}}\leftarrow\mathrm{S_{3}}\cup\mathrm{S_{2}}[k][q]$ $\triangleright$ save diverse sample
@@
43:return $\mathrm{S_{3}}$ $\triangleright$ final synthetic database for downstream training
@@ 3.5 Noisy Synthetic Data Filter
Although visually realistic and smooth (good cases, Fig. 6), sequence synthesis may still suffer from class semantics misalignment, cross-frame/slice inconsistency or over-consistency (i.e., almost static clip), and inter-clip similarity. For instance, in Fig. 6(b), the synthesized carotid clip category is wrong, which should be moderate rather than mild. In Fig. 6(f, h, j), the generated clip includes abrupt changes in anatomical structures. Hence, blindly using all synthetic clips for classifier learning will significantly cause a performance drop due to noisy samples. Our work proposes a noisy synthetic data filter to adaptively remove harmful generated clips at class semantics and sequential levels, as illustrated in Algorithm 1. This can also link the synthetic data to downstream tasks, thus potentially achieving a higher performance upper bound.
@@
Figure 6: Typical good and bad synthetic samples of five datasets using different banks of conditions, with representative failure patterns highlighted. Conditional controls include class label (C), text (T), image prior (I), and motion field (MF). See more synthetic results in Supplementary Material.
Class Semantics Misalignment Filtering. This scheme is designed to filter the generated clips whose visual contents mismatch the corresponding class labels. Motivated by anomaly detection (Kwon et al., 2020), we consider that such noisy synthetic samples will produce large losses if fed into a powerful classifier trained on real ones. Hence, a loss threshold can be set to identify and filter the anomalistic clips. Specifically, we first adopt a well-trained classifier in the real data domain to calculate case-wise cross-entropy loss values $\left\{l_{x}\right\}_{x=1}^{M}$ based on corresponding category $c$ for all synthetic clips in each group. Then, we set the average loss as the noise identification threshold of $c$. Last, we conduct class semantics misalignment filtering at the group level. If the loss of a synthetic clip with $c$ exceeds the threshold $L_{c}$, it will be viewed as a noisy sample and excluded from the synthetic pool. In our experiments, the classifier employed for semantic filtering is identical to the downstream classifier by default.
[…]
@@ 4 Experiments
We first benchmarked Ctrl-GenAug against 15 augmentation methods across 5 datasets with 4 medical modalities to comprehensively validate its effectiveness and generality. We further systematically assessed its downstream robustness using 11 popular classifiers trained under 3 paradigms. We then evaluated its performance in underrepresented high-risk sets and out-domain conditions. Finally, we conducted extensive ablation studies to examine the contributions of each component and the behavior of Ctrl-GenAug under different diffusion settings, backbone selections in the semantics misalignment filter, and varying annotation requirements.
Table 1: Experimental datasets and settings.
Carotid TUSC ACDC MosMedData MRNet
Modality US US Cine-MRI CT T2-weighted MRI
Organ Carotid artery Thyroid Heart Lung Knee
Patient 231 167 150 1,110 1,199
Clip 486 633 314 2,573 1,958
Mild/Moderate/Severe TI-RADS2-3/4/5 NOR/MINF/DCM/HCM/ARV Normal/Mild/Severe Negative/Positive
Class (train) 141/104/64 146/189/113 40/40/40/40/45 512/1,226/317 1,449/320
Class (test) 51/27/20 64/73/48 21/20/20/20/28 127/314/77 109/80
Class (OD) 37/24/18 - - - -
Sequence generator Training
Downsampling rate of VAE(8)   $\mathit{T}$(1,000)   AdamW(lr=1e-4)   lr scheduler(Cosine)
Warm(500 iterations)   Epochs(200)   Pretraining batch size(64)   Finetuning batch size(8)
Inference
Classifier-free guidance (Ho and Salimans, 2022) factor(7.5)   Sampling steps(200)
Classifiers Adam(default lr=1e-4, unless specified)   Batch size(default=8, unless specified)   lr scheduler(Cosine)
Epochs(100)   Standard augmentation practices(Random Color/Move/Gaussian/Rotation/Flip)
@@ 4.1 Datasets and Settings
MosMedData (Vladzymyrskyy et al., 2020) contains 3D lung CT sequences from 1,110 patients collected by municipal hospitals in Moscow. Each sequence was labeled according to the severity of lung tissue abnormalities with COVID-19, originally graded from CT-0 to CT-4 (CT-0: normal, CT-1: mild, CT-2/CT-3/CT-4: increasing severity). For analysis, CT-2, CT-3, and CT-4 were merged into a single severe category to alleviate class imbalance. We uniformly sampled the sequences to obtain 2,573 8-slice clips, of which 2,055 were used for training and 518 for testing.
MRNet (Bien et al., 2018) publicly releases 1,250 knee T2-weighted MRI exams, each comprising three orthogonal planes, from 1,199 patients acquired at the Stanford University Medical Center. Given the complexity of handling multiple imaging planes simultaneously, as suggested in (Khandelwal et al., 2018), we only included sagittal-plane sequences in the analysis for diagnosing anterior cruciate ligament (ACL) tears. Each sequence was then uniformly sampled into several 8-slice clips. The final dataset was split following the official protocol, resulting in 1,769 training clips and 189 testing clips.